Problem 5-14 (Static)

Here is a CPM network with activity times in weeks:

B(5) E(4) F(8)

a. Determine the critical path.

a. A-E-F-D

b. A-E-G-C-D

c. A-B-C-D


b. How many weeks will the project take to complete?


c. Suppose F could be shortened by two weeks and B by one week. By how many weeks would this reduce the project duration?
Answer #1

Answer: a. A-E-G-C-D

The given network paths are

A-E-F-D = 7+4+8+6 = 25 weeks

A-B-C-D = 7+5+6+6 = 24 weeks

A-E-G-C-D = 7+4+3+6+6 = 26 weeks

The critical path is A-E-G-C-D with the duration of 26 weeks

Answer b: The project takes 26 weeks to complete

This is based on the critical path of the project. The duration of critical path is 26 weeks

Answer c. 0 weeks

The activity B and F are not on the critical path and hence, it will not make any difference to the critical path.